# KIX: A Metacognitive Generalization Framework

### v1

@@ 3 Model
As a result, the agent follows a trajectory of activated interactions with objects and aids the formation of a task solution through meta-guided interactions. The proposed framework is generic to be applied to a wide range of domains.
4 Experiments
Experimental set up To assess the capabilities of the proposed agents, we require a challenging environment with ingredients for higher order reasoning to solve tasks. Our goal here is to investigate whether the introduction of meta-knowledge guided interactions benefits agents. MiniGrid world environments Chevalier-Boisvert et al. (2018) are suitable for our purpose. These are procedurally generated environments in which the agent interacts with objects such as doors, keys, balls, and boxes.
Figure 3: MiniGrid ObstructedMaze Full environment is an appropriate testbed because it is partially observable, requires complex reasoning, and new tasks can be created.
The environments are challenging due to partially observability, sparsity of rewards and random goal locations. The MiniGrid-ObstructedMaze-Full environment, figure 3, in particular, is a difficult environment with only a few claims for solving it. Due to the large state space size, stochasticity and partial observability, agents cannot memorize but must learn.
In the environment, the agent is initially placed in the middle room, with a 7x7 view. At the low level, the agent can move forward, left, right, pick, drop, and toggle, and its goal is to pick up the blue ball, which is randomly placed in one of the corner rooms. Some of the doors are locked and require a key of the same color to open. Keys are hidden inside boxes, but the agent cannot see inside the boxes without toggling. The original task is referred as Task 0. Furthermore, we identify the following additional task scenarios that were not seen during training but were presented to the agent during testing:
@@ 4 Experiments
  • Dynamic changes in the environment (Task 3) - when some event occurs, the environment changes. During training, the goal location remained unchanged. To simulate, as soon as the agent enters the room with the goal, relocate the goal to one of the neighboring rooms.
Agent training
Initialize model weights $\theta,\phi$ , parameters, $T$ for  $t=1$ to $T$ do
       // collect exp for meta exp size K
@@
             make instance graph $G_{I}\leftarrow\Phi(X)$ map $G_{I}$ to $G_{K}$, $G_{K}\leftarrow\Psi(G_{I})$ proposal $(e_{i},\lambda_{p})\leftarrow\Upsilon(G_{K},G_{I},\pi)$ // 1st variant
             interaction exp following $\mu_{\lambda_{p}}$ // 2nd variant
             reaching exp following $\mu_{\text{reach}}$ if success then
                   interaction exp following $\mu_{\lambda_{p}}$
             end if
            collect exp $\mathcal{D}_{\tau_{p}}$ and compute $A_{p}$
@@
      collect meta exp $\mathcal{D}_{\tau_{m}}$ and compute $A_{m}$ // Update policies, both variants
       update $\pi$ by optimizing $\mathcal{L}_{m}(\theta)$ for  $\lambda_{p}$ do
@@
      // 2nd variant only
       update $\mu_{\text{reach}}$ by optimizing $\mathcal{L}_{\text{reach}}(\phi_{\text{reach}})$
@@
Algorithm 1 KIX
@@
We devise two variants, KIX1 and KIX2. Both variants learn meta-policy and interaction policies. However the second variant additionally learns a reachability policy, so it interacts with an object once it has reached it. All meta, interactions, and reachability policies are implemented as actor-critic PPO policies. Algorithm 1 provides a high-level overview of the implementation. The input state embeddings for interaction and reachability actor-critic networks are constructed by a convolutional network with three Conv2D layers with 16, 32, and 64 output channels, a max pooling layer added after the first convolution layer, and exponential linear units used as activation functions. The meta-level recommends various objects for interaction, so an activation indicator for the object of interest is added to a new input channel, creating a four-channel input. Actors and critics are fully connected networks, each with one hidden layer. Meta-level policies are learned over type space knowledge graphs. In a type space graph, node object type information is encoded as an attribute vector. Edges in type space graph represent type-level relationships between nodes, such as "visible," "adjacent," "carrying," and "activated." A one-hot encoded edge attribute vector is used to encode these relationships. Embeddings for meta-level actor-critic are provided by two graph attention convolution layers, each with 16-dimensional embeddings and four attention heads. During message passing, these layers share weights between source and target. The final graph attention convolution layer is used for graph readout with a global add pool. Meta-level actors and critics are fully connected networks, with the critic learning a single value and the actor learning the probability of taking a meta-level action, such as ’pickup’, ’drop’, ’reveal’, ’open’, or ’open with key’. The meta-level actions serve as interaction goals. The environment rewards the agent with $r=1-(c/c_{T})$ where $c$ is the number of steps taken and $c_{T}$ is the max possible steps. For interaction policy learning, the agent receives a reward of 1 when it has a successful interaction. At the meta-level it is rewarded with 0.1 for a successful proposed interaction and $1+r$ for achieving the environment goal. A base agent is also implemented, with a similar network and a PPO policy, and it receives environmental rewards. Grid search is used to select hyperparameters. All the agents are trained for the same number of environment steps using 32 parallel workers.
Figure 4: The return profiles of agents performing various evaluation tasks show that the proposed agents have higher returns on all tasks and benefit from meta-guided interactions.
@@
Meta-guided interactions aid agents Our proposal is motivated by the idea that, rather than acting directly at the execution level, meta-level guided interactions with objects aid agents in generalizing while performing tasks. Because the agent’s environmental return is determined by the number of steps taken to complete the task, completing tasks in fewer steps yields a higher return. Therefore, the distribution of environmental returns is a reliable indicator of whether meta-level interaction recommendations benefit agents. We evaluate the distribution of environmental returns for each agent across all four tasks. For each task, the trained agents are run for 12K episodes by parallel workers, and their environmental rewards are recorded. In order to understand the effects of meta-guided interactions and to ensure a fair comparison between agents where differences in success rates are not a factor, we examine the top $k$ returns with $k=100$. Figure 4 illustrates the return profiles of agents. It is clear from the distribution of returns that the proposed agents have received a higher return on all tasks than the base agent, furthermore, this is even more so for dynamic tasks like task 3. It demonstrates that meta-guided interactions not only help agents complete their task goals in fewer steps, but also aid them to perform a wider range of tasks in static or dynamic environments. Furthermore, the higher returns of the second variant demonstrate that agents benefit from the introduction of reachability as a separate interaction. In the absence of a separate reachability policy, reaching is implicit in each interaction policy; however, separating allows for a shared learning of reachability followed by interactions which could be beneficial particularly in dynamically changing environments.
Figure 5: Wasserstein distances between traversal distributions, when agents complete tasks 1-3, and the trained task 0, provide comparative insights.
Agents profile While performing the tasks, agents interact with objects in the environment according to the recommended interaction policies; as a result, agents navigate and visit rooms. Because interactions are recommended by the meta-level, which is only trained on the training task, tracking agents’ room visits across all tasks provides an insight about their learning. Let $p_{0}$ be the distribution of agent room visits for trained task 0, and $p_{i},i\in\{1,2,3\}$ be the distributions for tasks 1–3. For both variants of the agent, we then compute the Wasserstein distance, a metric from optimal transport theory, between the visit distribution for tasks 1-3 and the trained task 0, as shown in figure 5. The distribution of visits to each agent is expected to be different from the training task, due to changes in the evaluation tasks. However, our interest is in how differently the two variants behave in evaluation tasks, as indicated by the computed metric. In task 1, the goal room remained unchanged, but a new relation was added to the evaluation. For both variants, the computed metric value for task 1 is similar, indicating that they transfer their learning to this task in a similar manner. However, there is a noticeable difference for tasks 2 and 3. In task 2, doors in the central room are locked with keys kept in the room - a static task, whereas in task 3, the goal is dynamically relocated based on observance - a dynamic task. The computed metric for the dynamic task for the second variant is smaller than the first variant, implying that its evaluation visit distribution is closer to that of the training task. It suggests that the learning in the second variant can better capture dynamic changes. When combined with the return profiles, it reinforces that having an additional reachability policy benefits the agent in a wider range of tasks, particularly dynamic tasks, and improves generalization.
@@ 5 Conclusion
In this work, we proposed a metacognitive framework that promotes generalist behaviors through three main contributions. First, generalist and flexible behaviors necessitate abstraction and composition, but current models lack natural abstractions and compositions that can preserve both object abstractions and relational knowledge structures. In our model, natural abstractions and compositions are represented and grounded using type space knowledge graphs which also preserve relational knowledge between objects. Second, rather than directly acting in the raw environment, interactions with objects play an important role in how an agent navigates the environment to complete tasks. To accomplish this, we propose an intermediary interaction level. The meta-level of the agent learns from type space knowledge graphs and recommends what objects to interact with, as well as what interaction should be. In this sense, our model also functions as an interaction goals generating machine. Third, interaction policies learned through interaction goals serve as a repository of policies for the agent to use in the execution space, providing for the necessary loose coupling for generalization via interactions as intermediaries. The agent’s meta-level recommends interactions based on knowledge structures and learned meta-policy, and the execution level applies interaction policies to complete tasks. We demonstrated that metacognitive knowledge guided interactions aid agents in generalization. Our framework is modular and adaptable, with the potential to have implications across domains through the induction or integration of knowledge structures, whether extrinsic, intrinsic, or derived, as well as control variables at the meta or interaction level.

### v2

@@ 3 Model
In summary, when viewing tasks as meta-policy guided interactions with objects, problems are naturally decomposed into interaction goals. Unlike the models that are tightly coupled with low-level representations through end-to-end policies, in the proposed framework, agent actions are aided by meta-level knowledge guided object interactions, providing a mechanism for generalist behaviors.
4 Results and Discussion
Experiments
Our goal is to examine whether meta-level knowledge-guided interactions benefit agents in task generalization. To assess the capabilities of the proposed agents, we need a challenging environment having ingredients for higher-order reasoning to solve tasks. MiniGrid world environments Chevalier-Boisvert et al. (2018) are suitable for this purpose; they are procedurally generated environments where the agent interacts with objects such as doors, keys, balls, and boxes.
ObstructedMaze-Full
ObstructedMaze-1Dlhb
Figure 3: Environments: ObstructedMaze-Full is used to experimentally evaluate the agent’s performance on the designed tasks, while ObstructedMaze-1Dlhb is used to evaluate the transferability of interaction policies.
The obstructed maze environments, shown in Figure 3, are particularly challenging due to partial observability, stochasticity, and sparse rewards, with only a few claims for solving it. In the MiniGrid-ObstructedMaze-Full environment, the agent is initially placed in the middle room, with a 7x7 view. At the low level, the agent can move forward, left, right, pick, drop, and toggle, and its goal is to pick up the blue ball, which is randomly placed in one of the corner rooms. Some of the doors are locked and require a key of the same color to open. Keys are hidden inside boxes, but the agent cannot see inside the boxes without toggling. The default task is referred to as Task 0. Furthermore, we identify the following new task scenarios, which were presented to the agent during testing but not during training:
@@ 4 Results and Discussion
  • Dynamic changes in the environment (Task 3) - when some event occurs, the environment changes. During training, the goal location remained unchanged. To simulate, as soon as the agent enters the room with the goal, relocate the goal to one of the neighboring rooms. It is only done once, not repeatedly.
Agents learning to interact using type space knowledge
initialize model weights $\theta,\phi$ , parameters, $T$
for  $t=1$ to $T$ do
       /* collect exp for meta exp size K */
@@
             observe a partial view $X$
             make instance graph $G_{I}\leftarrow\Phi(X)$
             map $G_{I}$ to $G_{K}$, $G_{K}\leftarrow\Psi(G_{I})$
             recommend $(e_{i},\lambda_{p})\leftarrow\Upsilon(G_{K},G_{I},\pi)$
             interact with object $e_{i}$ following $\mu_{\lambda_{p}}$
             collect interaction exps $\mathcal{D}_{\tau_{p}}$ and compute $A_{p}$
@@
      collect meta exp $\mathcal{D}_{\tau_{m}}$ and compute $A_{m}$
       /* Update policies */
       update $\pi$ by optimizing $\mathcal{L}_{m}(\theta)$
       for  $\lambda_{p}$ do
@@
Algorithm 1 KIX procedure
We devised two variants: KIX-A and KIX-R. Both variants learn meta-policy and a collection of interaction policies. While the KIX-A variant combines both reaching objects and interacting with them, the KIX-R variant learns an additional interaction policy referred to as the ‘reachability’ policy, which is used to reach objects prior to interacting with them. All policies, including meta and interactions, are implemented as actor-critic PPO policies. Algorithm 1 provides an overview of the procedure. The agent’s partially observed view $X$ contains objects with identifiers and attributes like color and type. The function $\Phi$ takes $X$ and creates an instance space graph by using visible objects as nodes, including the agent, and determining relationships between nodes such as ‘visible’, ‘adjacent’, and ‘carrying’. The ‘adjacent’ relation is derived using distance between the two objects. When the agent chooses an object to interact with, an ‘activated’ relation is added between the agent and the object node. The function $\Psi$ then constructs a type space graph from the instance space graph, encoding node types as attribute vectors. Edges in a type space graph have the same meaning as those in an instance space graph and are encoded as one-hot encoded edge attribute vectors.
Figure 4: Agents learning to interact: The learning progress of the agents KIX-A (blue) and KIX-R (red) is shown as returns over steps, which illustrate both the meta-level and interactions such as ‘open’, ‘open with key’, ‘reveal’, ‘pickup’ (not shown for KIX-R), ‘drop’, and ‘reach’ in the ObstructedMaze-Full environment.
@@
Agents learn how to perform low-level actions in the execution space using interaction policies. A convolutional network with three Conv2D layers with 16, 32, and 64 output channels, a max pooling layer added after the first convolution layer, and exponential linear units as activation functions is used to provide the input state embeddings for the interaction actor-critic. An activation indicator for the proposed interaction object is added as a new input channel, resulting in a four-channel input to the filters. Actor-critic networks are fully connected with a single hidden layer. Agents learn a meta-policy in the type space. Two graph attention layers with 4 heads and 16 dimensional embedding provide 64 dimensional input embeddings for the meta policy. During message passing, these layers share weights between source and target. The final graph attention layer is then used for graph readout, applying a global add pool. The meta-level actor and critic are both fully connected networks, with the critic learning a single value and the actor learning the likelihood of taking the meta-level actions ‘pickup’, ‘drop’, ‘reveal’, ‘open’, or ‘open with key’, each of which corresponds to an interaction policy.
Table 1: Training hyper-parameters
Parameter Value Parameter Value
meta exp size 8 gae lambda 0.95
exp size 64 entropy coeff 1e-04
meta minibatch size 64 value coeff 1.0
minibatch size 64 meta entropy coeff 1e-04
meta learning rate 1e-05 meta value coeff 1.0
learning rate 1e-05 ppo clip 0.2
gamma 0.99 ppo steps 4
The environment rewards the agents with $r=1-(c/c_{T})$ where $c$ is the number of steps taken and $c_{T}$ is the max possible steps. For interaction policies, the agent receives a reward of 1 when it has a successful interaction. At the meta-level, it is rewarded with 0.1 for a successful proposed interaction and $1+r$ for achieving the environment goal. The base agent, BASE, has a similar actor-critic architecture and receives environment rewards. The training hyper-parameters are given in Table 1. Learning rates were determined using grid search with values 1e-3, 1e-4, 1e-5, and 1e-6. All the agents are trained for $\sim 6.2\times 10^{8}$ steps, using 32 parallel workers.
As shown in Figure 4, as training progresses, agents learn to interact with objects and associate higher-level structured knowledge with these interactions using object types and knowledge graphs, as indicated by higher meta-level returns. This, in turn, leads to better interaction experiences and further improvements in interaction policies. Furthermore, the KIX-R variant first focuses on learning how to reach different objects, but once it learns to reach, it begins to learn interactions in fewer steps and with much higher returns compared to the KIX-A variant. It suggests that the agents learn higher-level relations and apply these learned relations and interactions with objects to solve problems.
Meta-level knowledge-guided interactions aid agents
@@
Our approach is motivated by the idea that meta-level guided interactions with objects help agents in generalizing while performing tasks, rather than acting directly at the execution level. Since the agent’s environmental return is determined by the number of steps taken to complete a task, completing tasks in fewer steps yields a higher return. Therefore, the distribution of environmental returns is a reliable indicator of the value of meta-level interaction recommendations for agents. We evaluate the distribution of environmental returns for each agent across four tasks. For each evaluation task, trained agents are run for 12K episodes by parallel workers, and their environmental rewards are tracked.
Figure 5: Evaluation return profiles of agents performing various tasks show that the proposed agents benefit from meta-level guided interactions.
To understand the effects of meta-guided interactions and ensure a fair comparison between agents, regardless of success rate differences, we examine the top $k$ returns with $k=100$. Figure 5 shows the return profiles of agents. It is evident from the distribution of returns that the proposed agents achieve higher returns on all tasks compared to the base agent, including dynamic tasks like Task 3. This demonstrates that meta-guided interactions not only assist agents in completing tasks in fewer steps, but also enable them to perform a wider range of tasks in both static and dynamic scenarios. Furthermore, higher returns of the second variant show that agents benefit from treating reachability as a separate interaction. In the absence of separate reachability, reaching is embedded into interactions. Separate reachability allows shared learning followed by interactions, which is particularly beneficial in dynamically shifting scenarios.
Transferability of policies
(a) Meta training
(b) Evaluation task returns
Figure 6: Evaluation return profiles show that interaction policies can be transferred from one environment to another.
The ObstructedMaze-1Dlhb environment is used to assess the transferability of interaction policies learned in one environment to another. In this environment, the agent’s goal is to pick up a blue ball hidden in a two-room maze. The rooms are separated by a door, which is obstructed by a ball and locked. The keys are hidden inside boxes. Agents employ previously acquired interaction policies from the base task in the ObstructedMaze-Full environment while only learning the meta-policies. During evaluation, agents use the learned meta-policy, but the interaction policies come from the ObstructedMaze-Full environment itself. Figure 6 shows the meta training returns and evaluation return profiles, similar to the main environment, for both agent variants. It is evident that agents have gained the ability to exploit interaction policies from the previous environment, and the second variant, which employs a separate reachability policy, appears to have benefited more.
@@ 5 Conclusion
In this work, we proposed a metacognitive framework that promotes generalist behaviors through three main contributions. First, generalist behaviors necessitate abstraction and composition; however, current models lack natural abstractions and compositions that can preserve both object abstractions and relational knowledge structures. In our model, type space knowledge graphs are used to represent and ground natural abstractions and compositions while also preserving relational knowledge between objects. Second, rather than directly acting in the environment, interactions with objects are critical in performing the tasks. To accomplish this, introduce an intermediary interaction level. The agent’s meta-level learns from type space knowledge graphs and makes recommendations on which objects to interact with and how to interact with them. In this sense, our model functions as an interaction goal-generating machine. Third, interactions employing interaction policies enable the necessary loose coupling required for generalization by acting as intermediaries. We demonstrated that metacognitive knowledge-guided interactions aid agents in generalization. Lastly, Our framework is modular and adaptable, allowing for the integration of knowledge structures and control variables at both the meta and interaction levels. This framework is promising, particularly for domains where higher-order knowledge and interaction compositions can be exploited, such as robotics, autonomous and cognitive systems.